OpenStudy (ttop0816):

it would be this one right?? Which is not an equation of the line that passes through the points (1, 1) and (5, 5)? y - x = 0 <--- x - y = 0 -x - y = 0

OpenStudy (anonymous):

No, it would be the last one. If you plug in the x and y values for each of the two points of the line, you can see that it does not satisfy the last equation. For example, plugging the first point into the last equation. -(1)-(1)=0 -2=0 -2 does not equal zero, therefore that point does not satisfy the equation, and the equation is not the equation of the line that passes through those points.